The correlation between the two stocks is 0.5.

Expected return (%)      Standard Deviation (%) Weight

Microsoft                 28                          42                    0.4                  

Coca-Cola              12.5                         21                    0.6

The standard deviation of rachels portfolio is 25.55%

James decides to follow her wife’s suggestion. But, he would prefer a lower volatility for his investment. He decides to invest 40% of his wealth into the T-bills, and 60% in Rachel’s portfolio.

What is the proportion of his investments in each asset?

He would invest ________% into the T-bills, ________% into Microsoft, and _________% into Coca-Cola.

Note: please fill in a whole number for each blank.
Answer #1

He would invest in:

T-bills = 40%

Microsoft = 0.4*60% = 24%

Coca-cola = 0.6*60% = 36%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.